OpenStudy (anonymous):

Find the value of X for which L is parallel to M 3x-43

OpenStudy (anonymous):



OpenStudy (anonymous):

you must make 3x-43=80 because of the alternate interier angles theorum. so solve for x 3x-43=80 +43 +43 3x=123 3x/3 = 123/3 x=41

OpenStudy (anonymous):

Thnx so much